OpenStudy (anonymous):

find the length of the diagonal of a 6 inch by 24 inch rectangle?

OpenStudy (goldphenoix):

 For this one, you must use the pythagoream theorm.

OpenStudy (anonymous):

ok I got everything but thesquare roo

OpenStudy (anonymous):

* square root

OpenStudy (anonymous):

is it 24.7 or leave it as radical 612

OpenStudy (anonymous):

oh no I got tht wrong my teacher came out with radical 232

OpenStudy (goldphenoix):

Okay, the pythagoream theorm is: \[\large \large a^2 + b^2 = c^2 \] C is the hypotenuse, or the longest side of a triangle.  \[\large \large 24^2 + 6^2 = c ^2 |= | 576 + 36 = c^2\] Next is to combine like term: 576 + 36 = 612. 612 = c^2 Next is to find the square root of 612. How can I do this? We can use a calculator, so the square root of it is ~24.74 inches. So the diagonal is 24.74 inches. I hope this help!